Please help me solve this problem!!!!!!!!!!
a bag contains rs 750 in the form of 1 rupee,50ps,25ps
coins in the form of 5:8:4.
Find the number of coins of each type

Answer by mananth(16946)

rupee,50ps,25ps
coins in the form of 5:8:4.

let the common multiple be x
rupee,50ps,25ps
coins in the form of 5x, 8x, 4x.
Total value of the coins
5x*1 + 0.5*8x+0.25*4x=750
5x+4x+x=750
10x=750
divide by 10
x=75
1 rupee coin 5*75 = 375 coins
50 p coins 8*75= 600 coins
25 paise coins 4*75=300 coins
Jai Hind
Lets check